# Navier-Stokes Equations with Fractional Dissipation and Associated Doubly Stochastic Yule Cascades

### v1

@@ 2 Navier-Stokes equations with Fractional Dissipation
If $\eta\in$① then $\xi\times\eta>0$. We have $\cos\theta_{\xi,\eta^{\perp}}=-\sin\theta_{\xi,\eta}$ and $\cos\theta_{\xi,\zeta^{\perp}}=\sin\theta_{\xi,\zeta}$. Then
@@
If $\eta\in$② then $\xi\times\eta<0$. We have $\cos\theta_{\xi,\eta^{\perp}}=\sin\theta_{\xi,\eta}$ and $\cos\theta_{\xi,\zeta^{\perp}}=-\sin\theta_{\xi,\zeta}$. Then
@@ 3 $(d,\gamma)$-DSY cascade
Let $\mathbb{T}=\{\theta\}\bigcup\left(\cup_{n=1}^{\infty}\{1,2\}^{n}\right)$ be the full binary tree rooted at $\theta$. The notion of doubly stochastic Yule (DSY) cascade was introduced in [part1_2021] as a tree-indexed family of random variables $\{Y_{v}=\lambda^{-1}_{v}T_{v}\}_{v\in\mathbb{T}}$, where $\{\lambda_{v}\}_{v\in\mathbb{T}}$ is a family of positive random variables independent of $\{T_{v}\}_{v\in\mathbb{T}}$, an i.i.d. mean-one exponentially distributed family of random variables.
@@
For $d\geq 1$ and $\gamma\in(\frac{1}{2},\frac{d+2}{4})$, a $(d,\gamma)$-DSY cascade is a tree-index family $\{Y_{v}=\lambda^{-1}_{v}T_{v}\}_{v\in\mathbb{T}}$ in which $\{T_{v}\}_{v\in\mathbb{T}}$, an i.i.d. mean-one exponentially distributed family of random variables and $\lambda_{v}=|W_{v}|^{2\gamma}$ where $\{W_{v}\}_{v\in\mathbb{T}}$ is a family of random vectors in $\mathbb{R}^{d}$ distributed as follows:
@@
    Given $W_{v}$, $T_{v}$ is exponentially distributed with intensity $|W_{v}|^{2\gamma}$, $W_{v1}$ is distributed as $H(\cdot|W_{v})$ and $W_{v2}=W_{v}-W_{v1}$,
@@
The DSY cascade associated with the fractional Navier-Stokes equations arises when recursively applying formula (2.14). Figure 2 is a cascade figure that illustrates the families $\{{Y}_{v}\}_{v\in\mathbb{T}}$ and $\{W_{v}\}_{v\in\mathbb{T}}$ induced from the recursion (2.14). Although the definition of the $\otimes$-product requires $d\geq 2$, the DSY cascade can be define for any $d\geq 1$.
@@
For each vertex $v\in\mathbb{T}$, we denote by $|v|$ the genealogical height of vertex $v$, and $v|j$ denotes the truncation up to the $j$’th generation with the convention that $v|0=\theta$. Regarding the time evolution of the $(d,\gamma)$ cascade $\{Y_{v}\}_{v\in\mathbb{T}}$, the recursion (2.14) terminates at a vertex $v\in\mathbb{T}$ if and only if
@@
The purpose of this section is to give criteria on explosion and non-explosion in terms of the dimension $d$ and the fractional power $\gamma$ (Section 3).
@@
Whenever $V(\xi,t)$ is a finite set, i.e. in the event $[S>t]$, one has
@@
On the nonexplosion event $[S>t]$, the solution process $\mathbf{X}$ is constructed over DSY $\mathbb{Y}$ via the recursion (2.14) as a $\odot$-product, in a suitable order, of the initial data $\chi_{0}$ evaluated on the $t$-leaves $\partial{V(\xi,t)}$ of the tree that “reached” time $t$. For example, in the tree realization in Figure 3,
@@
$\mathbb{E}[k(R)]={{c}_{d,\gamma}}\int_{{{\mathbb{R}}^{d}}}{k(|v|)\frac{1}{|\tilde{u}{{|}^{d+1-2\gamma}}|u-\tilde{u}{{|}^{d+1-2\gamma}}}d\tilde{u}}$. (3.7)
@@
With the aid of Mathematica (or using [gradshteyn, Formula 3.682, p.415]) to evaluate the above integral, one gets
@@
(see [part2_2021, Def. 2.5, Lem. 7.1]) The DSY $\{Y_{v}\}_{v\in\mathbb{T}}$ is said to be self-similar if for any initial state $W_{\emptyset}=\xi$ and for any $s\in\partial\mathbb{T}$, the ratios $\{R_{s|j}=|W_{s|j}|/|W_{s|j-1}|\}_{j\geq 1}$ are i.i.d. with a common distribution independent of $s$.
@@
By Section 3, the ratios $\{R_{v}\}_{v\in\mathbb{T}}$ are i.i.d. Thus, $\{Y_{v}\}_{v\in\mathbb{T}}$ is self-similar. The inverse ratio $\frac{1}{R}$ has a probability density function
@@
It is sufficient to consider the root vertex $v=\theta$. Let $W_{0}=\xi$, $W_{1}=\eta$, $\Phi_{1}=\phi_{1}$, $\Phi_{2}=\phi_{2}$, $u=\xi/|\xi|$, $\tilde{u}=\eta/|\xi|$. By the sine rule in the triangle made by $u$, $\tilde{u}$, $u-\tilde{u}$, we get
@@
In the classical case $d=3$ and $\gamma=1$, Le Jan and Sznitman [lejan, Prop. 1.2] remark that the angular vector $(\Phi_{1},\Phi_{2})$ is uniformly distributed in the triangle $\triangle$. It is easy to see that this is the case if and only if $d=3$ and $\gamma=1$.
@@
    $\mathbb{E}[R_{\max}^{-2\gamma}]=2{\bar{C}_{d,\gamma}}\int_{0}^{\pi/2}\int_{0}^{\phi_{1}}\frac{({{\sin}^{d+1-2\gamma}}({{\phi}_{1}}+{{\phi}_{2}})+{{\sin}^{d+1-2\gamma}}({{\phi}_{1}}-{{\phi}_{2}}))\sin^{2\gamma-2}\phi_{2}}{\sin^{2}\phi_{1}}d\phi_{2}d\phi_{1}$ (3.13)
    where $R_{\max}=\max\{R_{1},R_{2}\}$, $R_{v}=|W_{v}|/|W_{\theta}|$ for $v=1,2$.
@@
    If $\gamma\in[\frac{d+3}{6},\frac{d+2}{4})$ then $\mathbb{E}[R^{b}]>1$ for all $b>0$.
@@
    If $\gamma\in(\frac{1}{2},\frac{d+3}{6})$ then
    $\min\limits_{b>0}\mathbb{E}[R^{b}]=\mathbb{E}[R^{b^{*}}]={\bar{C}_{d,\gamma}}\int_{0}^{\pi}\int_{0}^{\pi-\phi_{1}}(\sin\phi_{2}\,\sin(\phi_{1}+\phi_{2}))^{\frac{d-1}{2}-\gamma}\sin^{2\gamma-2}\phi_{1}d\phi_{2}d\phi_{1}$
    where $b^{*}=\frac{d+3}{2}-3\gamma>0$.
@@
(i) Let $W_{0}=\xi$, $W_{1}=\eta$, $u=\xi/|\xi|$, $\tilde{u}=\eta/|\xi|$. On the event $\Phi_{1}\geq\Phi_{2}]$, we have $u~|R_{\max}$=|u-$\tilde${u}|. By (3.11),
$\mathbb{E}[R_{\max}^{-2\gamma}]$ $=$ $2\iint\limits_{{{\Delta}_{1}}}{|u-\tilde{u}|^{-2\gamma}{{S}_{d,\gamma}}({{\phi}_{1}},{{\phi}_{2}})d{{\phi}_{2}}d{{\phi}_{1}}}$
$=$ $2{\bar{C}_{d,\gamma}}\int_{0}^{\pi/2}{\int_{0}^{{{\phi}_{1}}}+\int_{\pi/2}^{\pi}\int_{0}^{\pi-\phi_{1}}{\frac{{{\sin}^{d+1-2\gamma}}({{\phi}_{1}}+{{\phi}_{2}})\sin^{2\gamma-2}\phi_{2}}{{{\sin}^{2}}{{\phi}_{1}}}d{{\phi}_{2}}d{{\phi}_{1}}}}$
where $\triangle_{1}=\{(\phi_{1},\phi_{2})\in\triangle:\phi_{1}\geq\phi_{2}\}$. Using the change of variables $\phi_{1}\mapsto\pi-\phi_{1}$, we can rewrite the above integral as (3.13).
(ii) Consider the case $\gamma\in[\frac{d+3}{6},\frac{d+2}{4})$. Denote
@@
$\mathbb{E}[R^{b}]=\int_{0}^{\infty}C_{d,\gamma}\frac{r^{2a-2\gamma-b}}{(r^{2}+1)^{a}}\tilde{F}_{d,\gamma}(r)dr=\mathbb{E}[R^{\bar{b}}]$
where $\bar{b}=2a-4\gamma+2-b<0$. The proof proceeds by noticing the logarithmic convexity of the moment generating function $\mathbb{E}[R^{x}]$ as follows. There exists $p\in(1,\infty)$ such that $\frac{b}{p}+\frac{\bar{b}}{p^{\prime}}=0$ where $\frac{1}{p}+\frac{1}{p^{\prime}}=1$. By Holder’s inequality,
@@
(iii) Consider $\gamma\in(\frac{1}{2},\frac{d+3}{6})$. By splitting $\int_{0}^{\infty}=\int_{0}^{1}+\int_{1}^{\infty}$ and using the change of variable $r\mapsto 1/r$ for the last integral, we get
By the Cauchy-Schwarz inequality, $r^{2\gamma-2+b}+r^{2a-2\gamma-b}\geq 2r^{a-1}$. Thus,
$\mathbb{E}[{{R}^{b}}]\geq\int_{0}^{1}{\frac{{{C}_{d,\gamma}}}{{{({{r}^{2}}+1)}^{a}}}2{{r}^{a-1}}{{{\tilde{F}}}_{d,\gamma}}(r)dr}=\mathbb{E}[{{R}^{b^{*}}}]$
where ${{b}^{*}}=a+1-2\gamma=\frac{d+3}{2}-3\gamma$. By (3.10), one can rewrite $\mathbb{E}[R^{b^{*}}]$ in terms of the angles as follows.

### v2

@@ 2 Navier-Stokes equations with Fractional Dissipation
If $\eta\in$① then $\xi\times\eta>0$. We have $\cos\theta_{\xi,\eta^{\perp}}=-\sin\theta_{\xi,\eta}$ and $\cos\theta_{\xi,\zeta^{\perp}}=\sin\theta_{\xi,\zeta}$. Thus
@@
If $\eta\in$② then $\xi\times\eta<0$. We have $\cos\theta_{\xi,\eta^{\perp}}=\sin\theta_{\xi,\eta}$ and $\cos\theta_{\xi,\zeta^{\perp}}=-\sin\theta_{\xi,\zeta}$. Thus
@@
For $\textup{(NSE)}_{d,\gamma}$ to be non-trivial, it is necessary that $d\geq 2$. Upon removal of the pressure and the divergence-free constraint, one arrives at the fractional Burgers equation, a well-known toy model for $\textup{(NSE)}_{d,\gamma}$, especially in the case $d=1$. Namely,
$u_{t}+(-\Delta)^{\gamma}u+\frac{1}{2}(u^{2})_{x}=0,\ \ u(\cdot,0)=u_{0},$ (2.20)
where $u=u(x,t)$, $x\in\mathbb{R}$, $t\geq 0$. For the corresponding range of values of $\gamma$, namely $\gamma\in(\frac{1}{2},\frac{3}{4})$, one can also interpret the function $\chi(\xi,t)=\frac{c_{0}\hat{u}(\xi,t)}{h(\xi)}$ probabilistically as the expected value of a solution process $\mathbf{X}$ defined in a similar manner as (2.14):
[…]
The fractional Navier-Stokes equations ($d\geq 2$) and the fractional Burgers equation ($d=1$) have essentially the same underlying stochastic structure called a doubly stochastic Yule (DSY) cascade, to be defined in the next section.
@@ 3 $(d,\gamma)$-DSY cascade
Let $\mathbb{T}=\{\emptyset\}\bigcup\left(\cup_{n=1}^{\infty}\{1,2\}^{n}\right)$ be the full binary tree rooted at $\emptyset$. The notion of a doubly stochastic Yule (DSY) cascade was introduced in [15] as a tree-indexed family of random variables $\{Y_{v}=\lambda^{-1}_{v}T_{v}\}_{v\in\mathbb{T}}$, where $\{\lambda_{v}\}_{v\in\mathbb{T}}$ is a family of positive random variables independent of the family of i.i.d. mean-one exponentially distributed random variables $\{T_{v}\}_{v\in\mathbb{T}}$.
@@
For $d\geq 1$ and $\gamma\in(\frac{1}{2},\frac{d+2}{4})$, a $(d,\gamma)$-DSY cascade is a tree-index family $\{Y_{v}=\lambda^{-1}_{v}T_{v}\}_{v\in\mathbb{T}}$ in which $\{T_{v}\}_{v\in\mathbb{T}}$ is an i.i.d. mean-one exponentially distributed family of random variables and $\lambda_{v}=|W_{v}|^{2\gamma}$, where $\{W_{v}\}_{v\in\mathbb{T}}$ is a family of random vectors in $\mathbb{R}^{d}$ distributed as follows:
@@
    Given $W_{v}$, $\lambda_{v}^{-1}T_{v}$ is exponentially distributed with intensity $|W_{v}|^{2\gamma}$, $W_{v1}$ is distributed as $H(\cdot|W_{v})$ and $W_{v2}=W_{v}-W_{v1}$,
@@
The DSY cascade associated with the fractional Navier-Stokes equations arises when recursively applying formula (2.14). Figure 2 is a cascade figure that illustrates the families $\{{Y}_{v}\}_{v\in\mathbb{T}}$ and $\{W_{v}\}_{v\in\mathbb{T}}$ induced from the recursion (2.14). Although the definition of the $\otimes$-product requires $d\geq 2$, the DSY cascade can be defined for any $d\geq 1$.
@@
For each vertex $v\in\mathbb{T}$, we denote by $|v|$ the genealogical height of vertex $v$, and $v|j$ denotes the truncation up to the $j$’th generation with the convention that $v|0=\theta$. Incorporating the time evolution of the $(d,\gamma)$ cascade $\{Y_{v}\}_{v\in\mathbb{T}}$, the recursion (2.14) terminates at a vertex $v\in\mathbb{T}$ if and only if
@@
The main purpose of this section is to give criteria on explosion and non-explosion in terms of the dimension $d$ and the fractional power $\gamma$ (Section 3.2).
@@
Whenever $V(\xi,t)$ is a finite set, i.e. on the event $[S>t]$, one has
@@
On the nonexplosion event $[S>t]$, the solution process $\mathbf{X}$ at (2.14) is a $\odot$-product, in a suitable order, of the initial data $\chi_{0}$ evaluated on the $t$-leaves $\partial{V(\xi,t)}$ of the tree that “reached” time $t$. For example, in the tree realization in Figure 3,
@@ 3.1 Distribution of the ratios
$\mathbb{E}[k(R)]={{c}_{d,\gamma}}\int_{{{\mathbb{R}}^{d}}}{k(|\tilde{u}|)\frac{1}{|\tilde{u}{{|}^{d+1-2\gamma}}|u-\tilde{u}{{|}^{d+1-2\gamma}}}d\tilde{u}}$. (3.7)
@@
With the aid of Mathematica (or using [25, Formula 3.682, p.415]) to evaluate the above integral, one obtains
@@
The proof that the sequence $\{R_{s|k}\}_{k\geq 1}$ is i.i.d. is a similar calculation by induction, conditioning and using the Markov property along rays. The details are left to the reader. ∎
@@
(see [16, Def. 2.5, Lem. 7.1]) The DSY $\{Y_{v}\}_{v\in\mathbb{T}}$ is said to be self-similar if for any initial state $W_{\emptyset}=\xi$ and for any $s\in\partial\mathbb{T}$, the ratios $\{R_{s|j}=|W_{s|j}|/|W_{s|j-1}|\}_{j\geq 1}$ are i.i.d. with a common distribution independent of $s$.
@@
By Section 3.1, the ratios $\{R_{s|k}\}_{k\geq 1}$ are i.i.d. Thus, $\{Y_{v}\}_{v\in\mathbb{T}}$ is self-similar. The inverse ratio $\frac{1}{R}$ has a probability density function
@@
It is sufficient to consider the root vertex $v=\theta$. Let $W_{0}=\xi$, $W_{1}=\eta$, $\Phi_{1}=\phi_{1}$, $\Phi_{2}=\phi_{2}$, $u=\xi/|\xi|$, $\tilde{u}=\eta/|\xi|$. By the law of sines in the triangle made by $u$, $\tilde{u}$, $u-\tilde{u}$, we get
@@
In the classical case $d=3$ and $\gamma=1$, Le Jan and Sznitman [28, Prop. 1.2] remark that the angular vector $(\Phi_{1},\Phi_{2})$ is uniformly distributed in the triangle $\triangle$. It is easy to see that this is the case if and only if $d=3$ and $\gamma=1$.
Remark 3.9.
The symmetry with respect the multiplicative group $(0,\infty)$ of the ratio $R$ noted in Section 3.1 is a consequence of the invariance of the joint distribution of $(R_{1},R_{2})$ with respect to the Kelvin inversion with respect to the unit sphere. This can proved using the joint distribution of the angles given in Section 3.1 as follows. Note that in terms of the angles the Kelvin inversion is the transformation (see Figure 4)
$(\phi_{1},\phi_{2})\rightarrow(\phi_{1},\phi_{3}),\qquad\phi_{3}=\pi-(\phi_{1}+\phi_{2}).$ (3.13)
Since $\sin(\phi_{1}+\phi_{2})=\sin(\phi_{3})$, the joint density of the angles is invariant iff
$d+1-4\gamma=2\gamma-2$
Figure 4: Kelvin inversion $\eta\mapsto\eta^{*}$, where $W_{1}=\eta$ and $\eta^{*}=\frac{1}{|\eta|^{2}}\eta$.
i.e. iff $\gamma=\tfrac{d+3}{6}$. The coordinate transformation (3.13) will be used in the proof of Section 3.2, highlighting the importance of this symmetry in the explosion properties of the DSY cascade $\mathbb{Y}$.
Remark 3.10.
Alternatively, instead of angles $\Phi_{1}$ and $\Phi_{2}$, one can describe the basic triangle formed by $W_{v}$, $W_{v1}$, and $W_{v2}$ in terms of (normalized) side lengths: $1=|W_{v}|/|W_{v}|$, $R_{1}=|W_{v1}|/|W_{v}|$, and $R_{2}=|W_{v2}|/|W_{v}|$, with the joint distribution of $R_{1}$ and $R_{2}$ given by the pdf
$h_{d,\gamma}(r_{1},r_{2})=\tilde{C}_{d,\gamma}\frac{A(1,r_{1},r_{2})^{d-3}}{(r_{1}r_{2})^{d-2\gamma}}{\mathds{1}_{D}}(r_{1},r_{2}),$ (3.14)
where, using Heron’s triangle formula for area,
$A(1,r_{1},r_{2})=\frac{1}{4}\sqrt{(1+r_{1}+r_{2})(1+r_{1}-r_{2})(1+r_{2}-r_{1})(r_{1}+r_{2}-1)}$
is the area of the triangle with sides $1$, $r_{1}$, $r_{2}$. Here,
$D=\{(r_{1},r_{2})\in\mathbb{R}^{2}:r_{1},r_{2}>0,r_{1}+r_{2}>1,|r_{1}-r_{2}|<1\}.$
As we will see, moments of random variables $R_{v}$ will play a defining role in the explosive/non-explosive nature of the corresponding DSY cascades.
@@
    For all $b\in\mathbb{R}$, $\mathbb{E}[R^{b}]=\mathbb{E}[R^{\bar{b}}]$ where $\bar{b}=d+3-6\gamma-b$.
@@
    Let $b^{*}=\frac{d+3}{2}-3\gamma$. Then
    $\min\limits_{b\in\mathbb{R}}\mathbb{E}[R^{b}]=\mathbb{E}[R^{b^{*}}]={\bar{C}_{d,\gamma}}\int_{0}^{\pi}\int_{0}^{\pi-\phi_{1}}(\sin\phi_{2}\,\sin(\phi_{1}+\phi_{2}))^{\frac{d-1}{2}-\gamma}\sin^{2\gamma-2}\phi_{1}d\phi_{2}d\phi_{1}.$
@@
    $\mathbb{E}[R^{b}]\geq 1$ for all $(b,\gamma)\in(0,\infty)\times[\frac{d+3}{6},\frac{d+2}{4})\bigcup\,(-\infty,0)\times(\frac{1}{2},\frac{d+3}{6}]$.
@@
(i) Let $a=\frac{d+1}{2}-\gamma>0$ and
@@
$\mathbb{E}[R^{b}]=\int_{0}^{\infty}C_{d,\gamma}\frac{r^{2a-2\gamma-b}}{(r^{2}+1)^{a}}\tilde{F}_{d,\gamma}(r)dr=\mathbb{E}[R^{\bar{b}}].$
(ii) By splitting $\int_{0}^{\infty}=\int_{0}^{1}+\int_{1}^{\infty}$ and using the change of variable $r\mapsto 1/r$ for the last integral, we get
By the Cauchy-Schwarz inequality, or since the geometric mean cannot exceed the arithmetic mean, $r^{2\gamma-2+b}+r^{2a-2\gamma-b}\geq 2r^{a-1}$. Thus,
$\mathbb{E}[{{R}^{b}}]\geq\int_{0}^{1}{\frac{{{C}_{d,\gamma}}}{{{({{r}^{2}}+1)}^{a}}}2{{r}^{a-1}}{{{\tilde{F}}}_{d,\gamma}}(r)dr}=\mathbb{E}[{{R}^{b^{*}}}].$ (3.16)
By (3.10), one can rewrite $\mathbb{E}[R^{b^{*}}]$ in terms of the angles as
(iii) In the case $(b,\gamma)\in(0,\infty)\times[\frac{d+3}{6},\frac{d+2}{4})\bigcup\,(-\infty,0)\times(\frac{1}{2},\frac{d+3}{6}]$, note that $b$ and $\bar{b}$ have different signs. There exists $p\in(1,\infty)$ such that $\frac{b}{p}+\frac{\bar{b}}{p^{\prime}}=0$ where $\frac{1}{p}+\frac{1}{p^{\prime}}=1$. The proof proceeds by noticing the logarithmic convexity of the moment function $x\mapsto\mathbb{E}[R^{x}]$ as follows. By Holder’s inequality,
@@
3.2 Explosion and hyper-explosion of $(d,\gamma)$-DSY cascades
We start by recalling the explosion and non-explosion criteria for self-similar $(d,\gamma)$-DSY cascades from [16]. Note that part (b) in the theorem below is in fact an improvement of [16, Prop. 2.7], which can be applied to general self-similar DSY considered there. We also establish two new conditions for hyperexplosion and nonhyperexplosion, which for simplicity are formulated in the context of the $(d,\gamma)$-DSY, but transfer in a straightforward way to general self-similar DSY considered in [16].
@@ 3.2 Explosion and hyper-explosion of $(d,\gamma)$-DSY cascades
A self-similar $(d,\gamma)$-DSY cascade is
  1. (a)
    non-explosive if $\mathbb{E}[R^{b}]<1/2$ for some $b>0$;
  2. (b)
    explosive if $\mathbb{E}[\ln R_{\max}]>0$, where $R_{\max}=\max\{R_{1},R_{2}\}$, $R_{1}=\frac{|W_{1}|}{|W_{\theta}|}$, $R_{2}=\frac{|W_{2}|}{|W_{\theta}|}$.
  3. (c)
    If $\mathbb{E}[\ln R_{\min}]<0$, where $R_{\min}=\min\{R_{1},R_{2}\}$ then the cascade is a.s. not hyperexplosive.
  4. (d)